The bound electron in a hydrogen atom starts off in an excited state and releases a photon with a wavelength of 95.0 nm as it transitions to the ground state.

What orbital (n) did the electron start in?
1
Expert's answer
2020-05-29T09:59:18-0400

Wavelength of photon is "\lambda=95nm" .

Let us consider the electron was in "n=n_1" orbital and transits to "n=1".


Now, from Bohr's principle of hydrogen atom,we know that the amount of energy released by electron in transition is equal to energy of the emitted photon, thus we get

"\\Delta=\\frac{hc}{\\lambda}=13.6\\bigg(\\frac{1}{1^2}-\\frac{1}{n_1^2}\\bigg)eV\\\\\n\\implies \\frac{1240eV\\cdot nm}{\\lambda}=13.6\\bigg(\\frac{1}{1^2}-\\frac{1}{n_1^2}\\bigg)eV\\\\\n\\implies \\frac{1240eV\\cdot nm}{95nm}=13.6\\bigg(\\frac{1}{1^2}-\\frac{1}{n_1^2}\\bigg)eV\\\\\n\\implies \\frac{1}{n_1^2}=1-0.9597=0.0403\u202c\\\\\\implies n_1=\\sqrt{\\frac{1}{0.0403\u202c}}=4.98"

But,"n" is always integer hence,"n_1=5"